OVHCloud Web Hosting Status

Current status
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#5483 — Internal routing, collect operator
Scheduled Maintenance Report for Web Cloud
Since February, there has been an internal routing problem at a
collect operator. Instead of moving directly to our links dedicated
to the collect, the traffic moved through our peering. The peering being on Paris,
all the traffic moves through Paris even if you had chosen an LNS at
Bordeaux, Lyon or Marseille.

The routing has been rectified and you can now go directly through the chosen
LNS and you are faster on the OVH network.


Date: 2011-12-15 13:43:23 UTC
Everything is OK from OVH and sfr's side.

The switch is done.

Date: 2011-12-09 01:01:56 UTC
Paris done.

All PPP sessions are now on the new architecture.

Tomorrow we will check with SFR that everything is alright from both sides.

Date: 2011-12-09 01:01:46 UTC
Lyon done.

Date: 2011-12-09 01:01:29 UTC
Bordeaux done.

Date: 2011-12-09 01:01:12 UTC
LNS Marseille done.

Date: 2011-12-09 01:00:57 UTC
LNS Roubaix done.

Date: 2011-12-09 01:00:40 UTC
The first accounts on roubaix are cut.
The cuts takes only a few seconds.

Date: 2011-12-09 00:59:30 UTC
The migration from SFR's side was done.

Now we are going through the concentration BAS located in Paris. The connections are directly established from the DSLAM if it is configured on Level 3 or by BAS as close as possible if the DSLAM are on Level 2.

The old sessions are still connected by the BAS in Paris. For the next reconnection, the new path will be taken.

Tonight we will resetet all PPP connections in order to switch to the new architecture.

Date: 2011-12-09 00:52:06 UTC
We modified the network characteristics of the links between OVH and SFR in order to make our tests.

The LNS Paris / Roubaix / Marseille / Bordeaux / Lyon are ready.

Date: 2011-12-09 00:50:31 UTC
We will make some tests on the links of tests in order to validate the architecture.

Date: 2011-11-03 12:53:10 UTC
There are problems persisting on the routing between the collect operator and OVH. The incident is being resolved between our teams.

Date: 2011-05-31 18:43:41 UTC
Some problems persist on the routing between the collecting operator
and OVH. The incident is being solved between our teams.
Posted May 30, 2011 - 10:49 UTC